OpenStudy (anonymous):

Help what is the inverse of: y = (e^x)/(1+7e^x) I got: y = (e^x)/(1+7e^x) x = (e^y) / (1 + 7e^y) x + 7xe^y = e^y x = e^y - 7xe^y x = (e^y)(1 - 7x) (x)/(1 - 7x) = e^y How would I bring down the y? Thank you.

OpenStudy (anonymous):

take the natural log (ln) of both sides \[\ln (x/(1-7x)) = \ln(e ^{y})\] Due to the properties of logs, the y can be dropped, and ln(e) = 1 Therefore: y = ln(x/(1-7x))

OpenStudy (anonymous):

I see ok thanks. ill try it out.

OpenStudy (anonymous):

Thanks. it workwed